-2x + 0y = -14
x + 11y = 18
Solve using the substitution method. Show your work. If the system has no solution or an infinite number of solutions, state this.

Answer by mananth(16946)

-2x + 0y = -14
-2x=-14
/-2
x=7
x + 11y = 18
plug the value of x
7+11y=18
11y=18-7
11y=11
/11
y=1
..
there is only one solution x=7,y=1
...
m.ananth@hotmail.ca


Solve using the substitution method. Show your work. If the system has no solution or an infinite number of solutions, state this.